# Xingwang Li, , Dusit Niyato, and Chau Yuen

### v1

@@ III-B Kernel Ridge Regression-Based Channel Estimation
While KNN provides a simple heuristic for interpolation, it lacks a strong theoretical foundation for optimality. Kernel-based methods, on the other hand, offer a powerful and principled framework for non-linear function approximation, rooted in the theory of reproducing kernel Hilbert spaces (RKHS) [28]. A prominent and practical example of this approach is kernel ridge regression (KRR) [27]. The KRR models the unknown channel function $\mathbf{h}(\bm{\zeta})$ as an element of the RKHS, where the estimation for any target shape $\mathbf{\bm{\zeta}}_{\text{target}}$ is represented as a linear combination of kernel functions evaluated at the pilot locations. The estimate is given by
@@
where $k(\cdot,\cdot)$ is a symmetric and positive-definite kernel function that measures the similarity between two shape vectors, and $\mathbf{c}_{m}\in\mathbb{C}^{N}$ is the coefficient vector to be determined.
@@
$k(\bm{\zeta}_{a},\bm{\zeta}_{b})=\exp\left(-\gamma\|\bm{\zeta}_{a}-\bm{\zeta}_{b}\|^{2}\right),$ (23)
where $\gamma>0$ is a hyperparameter that controls the bandwidth or length-scale of kernel. Parameters $a$ and $b$ denote different FIM deformation shapes.
@@
where $\mathbf{H}_{\text{pilots}}=[\hat{\mathbf{h}}(\bm{\zeta}_{1}),\dots,\hat{\mathbf{h}}(\bm{\zeta}_{M})]^{T}\in\mathbb{C}^{\times N}$ is the matrix of all pilot channel measurements. $\mathbf{K}\in\mathbb{R}^{M\times M}$ is the kernel matrix with entries $[\mathbf{K}]_{ij}=k(\bm{\zeta}_{i},\bm{\zeta}_{j})$. $\lambda\geq 0$ is the regularization parameter. $\|\cdot\|_{F}$ denotes the Frobenius norm.
@@
Finally, to estimate the channel at a new target shape $\mathbf{\bm{\zeta}}_{\text{target}}$, we first compute the $M\times 1$ vector of kernel similarities, $\mathbf{k}_{\text{target}}$, where $[\mathbf{k}_{\text{target}}]_{m}=k(\mathbf{\bm{\zeta}}_{\text{target}},\bm{\zeta}_{m})$. The final channel estimate is then synthesized as
@@
Compared to KNN, KRR finds an optimal set of weights by solving a global optimization problem. KRR provides a numerically stable and computationally efficient method, requiring only the inversion of a small and well-behaved $M\times M$ kernel matrix. However, the interpolation-based channel estimation method does not leverage prior information about the FIM channel.
@@ III-C Sparsity-Based Parametric Channel Estimation
The goal is to recover the sparse vector $\mathbf{x}$ from the low-dimensional measurements $\{\hat{\mathbf{h}}(\bm{\zeta}_{m})\}_{m=1}^{M}$ with the measurement noise $\mathbf{n}$. We can stack the measurements into a single system of equations:
$\begin{split}\hat{\mathbf{h}}_{\text{stack}}=\begin{bmatrix}\hat{\mathbf{h}}(\mathbf{\bm{\zeta}}_{1})\\ \vdots\\ \hat{\mathbf{h}}(\mathbf{\bm{\zeta}}_{M})\end{bmatrix}=\begin{bmatrix}\mathbf{A}(\mathbf{\bm{\zeta}}_{1})\\ \vdots\\ \mathbf{A}(\mathbf{\bm{\zeta}}_{M})\end{bmatrix}\mathbf{x}+\mathbf{n}.\end{split}$ (28)
This is a standard compressed sensing problem, which can be solved using various sparse recovery algorithms. Orthogonal matching pursuit (OMP) is an efficient greedy algorithm for this sparse channel estimation task [30], which iteratively identifies the support set of the sparse vector $\mathbf{x}$, i.e., the indices of the non-zero entries. Let $\hat{\mathbf{x}}$ be the recovered sparse parameter vector, the channel for any continuous shape $\mathbf{\bm{\zeta}}_{\text{target}}$ can be synthesized as $\hat{\mathbf{h}}_{\text{OMP}}(\mathbf{\bm{\zeta}}_{\text{target}})=\mathbf{A}(\mathbf{\bm{\zeta}}_{\text{target}})\hat{\mathbf{x}}$. This approach elegantly solves both the low-overhead and continuous-space estimation challenges by moving from direct channel interpolation to physical model reconstruction.
The performance of OMP is theoretically guaranteed by the properties of the sensing dictionary $\mathbf{\Phi}\in\mathbb{C}^{MN\times D}$, which is constructed from the pilot shapes $\{\bm{\zeta}_{m}\}$ and a grid of potential angles. A crucial metric is the mutual coherence $\mu(\mathbf{\Phi})$, which measures the maximum similarity between any two distinct atoms in the dictionary:
@@
A low mutual coherence is a sufficient condition for OMP to guarantee exact recovery of the sparse signal in the noiseless case, and it ensures stability and robustness in the presence of noise. Since the sensing dictionary $\mathbf{\Phi}$ is constructed from the pilot shapes $\{\bm{\zeta}_{m}\}$, we can optimize these shapes to minimize the mutual coherence, thereby creating the best possible conditions for OMP. The optimization problem is given by
@@
To find a high-quality solution to the coherence minimization problem without exhaustive search, we propose a constructive greedy algorithm that builds the set of optimal pilot shapes sequentially. At each step, it selects a new shape from a large pool of candidates that, when added to the already selected set, results in the minimal possible mutual coherence for the newly formed dictionary.
@@
The above conventional model-based channel estimation frameworks, while foundational, exhibit certain inherent limitations when confronted with the high complexity of the FIM channel estimation problem. The interpolation-based approach, provides the stable linear estimate, yet the fundamental relationship between the FIM deformation vector $\bm{\zeta}$ and the resulting channel vector $\mathbf{h}(\bm{\zeta})$ is highly non-linear. Consequently, the linear estimator or kernel-based models are inherently suboptimal and cannot fully capture the intricate channel dynamics. On the other hand, parametric methods based on sparse recovery are constrained by rigid assumptions that may not hold in practice. Their efficacy critically hinges on two conditions: 1) the channel must be sufficiently sparse, which is often not the case in rich scattering environments, and 2) the true path angles must align with a predefined discrete grid $\mathbf{A}(\bm{\zeta})$, making the methods highly susceptible to basis mismatch errors. These limitations motivate the development of a more flexible and data-driven approach.
@@
IV Learning-Based Chanel Estimation with Fourier Neural Operator for FIM Systems
@@
To overcome the aforementioned limitations of model-based channel estimation approaches, in this section, we develop a learning-based framework that moves beyond explicit model assumptions of linearity or sparsity. Specifically, we first provide a theoretical justification for learning-based channel estimation from a Bayesian inference perspective. Then, we reframe the channel estimation problem as one of learning a continuous, non-linear operator that maps any given FIM deformation shape to its corresponding channel response. For this purpose, we propose an H-FNO architecture uniquely suited for this task due to its inherent mesh-independence and its design for learning solutions to physical systems governed by partial differential equations (PDEs), such as the wave equation that underpins wireless propagation.
@@ IV-A FIM Channel Estimation from a Bayesian Perspective
In essence, channel estimation can be framed as a Bayesian problem of inferring the most plausible channel function $\bar{h}$, given a set of noisy measurements, $\{\hat{\mathbf{h}}(\bm{\zeta}_{m})\}_{m=1}^{M}$. According to Bayes’ theorem, the posterior probability distribution of the channel function is given by [31]
$p\left(\bar{h}\mid\{\hat{\mathbf{h}}(\bm{\zeta}_{m})\}\right)\propto p\left(\{\hat{\mathbf{h}}(\bm{\zeta}_{m})\}\mid\bar{h}\right)p(\bar{h}),$ (31)
where the likelihood term $p(\{\hat{\mathbf{h}}(\bm{\zeta}_{m})\}\mid\bar{h})$ is determined by the noise model. Based on the AWGN model in Eq. (II-C), the likelihood of the measurements is a Gaussian distribution centered around the true channel vectors $\bar{h}(\bm{\zeta}_{m})$. For independent measurements, this yields
$p\left(\{\hat{\mathbf{h}}(\bm{\zeta}_{m})\}\mid\bar{h}\right)\propto\exp\left(-\frac{1}{2\sigma_{w}^{2}}\sum_{m=1}^{M}\left\|\hat{\mathbf{h}}(\bm{\zeta}_{m})-\bar{h}(\bm{\zeta}_{m})\right\|_{2}^{2}\right).$ (32)
@@
The goal of channel estimation is to find $\bar{h}$, which corresponds to the Maximum A Posteriori (MAP) estimate. The MAP estimator maximizes the posterior probability, which is equivalent to minimizing the negative log-posterior:
$\bar{h}_{\text{MAP}}=\arg\min_{\bar{h}}\left(-\log p\left(\{\hat{\mathbf{h}}(\bm{\zeta}_{m})\}\mid\bar{h}\right)-\log p(\bar{h})\right).$ (33)
By substituting the Gaussian likelihood, the negative log-likelihood term becomes equivalent to a squared $\ell_{2}$-norm, which acts as a data fidelity term. This transforms the MAP estimation problem into the canonical form of a regularized optimization problem [29]:
@@
$\begin{split}\bar{h}_{\text{MAP}}=\arg\min_{\bar{h}}\left(\sum_{m=1}^{M}\left\|\hat{\mathbf{h}}(\bm{\zeta}_{m})-\bar{h}(\bm{\zeta}_{m})\right\|_{2}^{2}-2\sigma_{w}^{2}\log p(\bar{h})\right).\end{split}$ (34)
@@
This final expression provides that any channel estimator can be interpreted as a solver for this problem. The first term enforces fidelity to the measurements, while the second term $-2\sigma_{w}^{2}\log p(\bar{h})$, acts as a regularizer that is fundamentally defined by the choice of the prior $p(\bar{h})$. Consequently, the choice of the estimator is fundamentally governed by the choice of the prior $p(\bar{h})$, which represents our knowledge or assumptions about the channel structure. The model-based methods in Section III make explicit, but often restrictive, prior assumptions, e.g., a strict sparsity prior for OMP estimator. To motivate a more general and physically-grounded prior that can serve as the foundation for a learning-based approach, we now analyze the inherent spectral properties of the FIM channel model itself. As defined in Eq. (4), the FIM channel is a linear superposition of $S=L\times G$ steering vectors, each corresponding to a distinct path angle $(\theta_{s},\phi_{s})$. The response of the $n$-th antenna to the $s$-th path can be expressed as a complex exponential of its spatial coordinates $(x_{n},z_{n})$:
$\left[\mathbf{a}(\bm{\zeta},\theta_{s},\phi_{s})\right]_{n}=e^{j\frac{2\pi}{\lambda}(y_{n}\sin\theta_{s}\sin\phi_{s})}\cdot e^{j\left(k_{x,s}x_{n}+k_{z,s}z_{n}\right)},$ (35)
@@
where $\alpha_{k}^{\prime}(\bm{\zeta})$ is the effective complex gain of the $k$-th path, which incorporates the original path gain $\alpha_{k}$ and the spatially-varying phase shift induced by the FIM deformation.
Since the Fourier transform of a complex sinusoid is a Dirac delta function, the spatial spectrum $\tilde{\mathbf{h}}(\bm{\zeta};\omega_{x},\omega_{z})$ of the channel is non-zero only at a discrete set of $S$ spatial frequencies. This derivation proves a fundamental property: for any given deformation $\bm{\zeta}$, the FIM channel is inherently sparse in the spatial frequency domain, which provides a rigorous physical justification for selecting an superior prior $p(\bar{h})$. A function that is sparse in the frequency domain is a specific instance of a broader class of functions whose Fourier coefficients decay rapidly. Consequently, a powerful and general prior for this problem should favor functions with this spectral decay characteristic. By operating directly in the Fourier domain and inherently focusing on a limited number of modes [32], the FNO architecture is naturally aligned with the physical properties of the FIM channel.

### v2

@@ III-B Kernel Ridge Regression-Based Channel Estimation
While KNN provides a simple heuristic for interpolation, it lacks a strong theoretical foundation for optimality. Kernel-based methods offer a powerful and principled framework for non-linear function approximation, rooted in the theory of reproducing kernel Hilbert spaces (RKHS) [36]. A prominent and practical example of this approach is kernel ridge regression (KRR) [35]. The KRR models the unknown channel function $\mathbf{h}(\bm{\zeta})$ as an element of the RKHS, where the estimation for any target shape $\mathbf{\bm{\zeta}}_{\text{target}}$ is represented as a linear combination of kernel functions evaluated at the pilot locations. The estimate is given by
@@
where $k(\cdot,\cdot)$ is a symmetric and positive-definite kernel function that measures the similarity between two shape vectors, and $\mathbf{c}_{m}\in\mathbb{C}^{N}$ is the coefficient vector to be determined. The set of pilot shapes $\bm{\zeta}_{m}$ is also generated from the 2D Fourier basis, in which the systematic and non-redundant sampling of the deformation space provided by the Fourier basis helps to form a well-conditioned kernel matrix.
@@
$k(\bm{\zeta}_{a},\bm{\zeta}_{b})=e^{\left(-\gamma\|\bm{\zeta}_{a}-\bm{\zeta}_{b}\|^{2}\right)},$ (21)
where $a$ and $b$ denote different FIM deformation shapes. The hyperparameter $\gamma>0$ controls the bandwidth or length-scale of kernel. Specifically, $\gamma$ encodes a prior assumption about the smoothness of the channel function. A small $\gamma$ implies a long length-scale, assuming the channel varies smoothly with deformation, while a large $\gamma$ implies a short length-scale, allowing the function to change more rapidly.
@@
where $\mathbf{H}_{\text{pilots}}=[\hat{\mathbf{h}}(\bm{\zeta}_{1}),\dots,\hat{\mathbf{h}}(\bm{\zeta}_{M})]^{T}\in\mathbb{C}^{M\times N}$ is the matrix of all pilot channel measurements. $\mathbf{K}\in\mathbb{R}^{M\times M}$ is the kernel matrix with entries $\mathbf{K}]_{ij}=k(\bm{\zeta}_{i},\bm{\zeta}_{j})$. The regularization parameter $\lambda\geq 0$ controls the trade-off between data fidelity and the smoothness prior set. A larger $\lambda$ imposes a stronger penalty on model complexity, enforcing a smoother solution that prevents overfitting to noisy pilot measurements. The optimal value of $\lambda$ cannot be determined analytically. In this work, it is selected using standard model selection techniques by cross-validation.
@@
Finally, to estimate the channel at a new target shape $\mathbf{\bm{\zeta}}_{\text{target}}$, we first compute the $M\times 1$ vector of kernel similarities, $\mathbf{k}_{\text{target}}$, where $\mathbf{k}_{\text{target}}]_{m}=k(\mathbf{\bm{\zeta}}_{\text{target}},\bm{\zeta}_{m})$. The final channel estimate is then synthesized as
@@
Compared to KNN, KRR finds an optimal set of weights by solving a global optimization problem. KRR provides a numerically stable and computationally efficient method, requiring only the inversion of a small and well-behaved $M\times M$ kernel matrix. However, the interpolation-based channel estimation method does not leverage prior information about the channel.
@@ III-C Sparsity-Based Parametric Channel Estimation
The goal is to recover the sparse vector $\mathbf{x}$ from the low-dimensional measurements $\{\hat{\mathbf{h}}(\bm{\zeta}_{m})\}_{m=1}^{M}$ with the measurement noise $\mathbf{n}\sim\mathcal{CN}(0,\sigma^{2}\mathbf{I})$. We can stack the measurements into a single system of equations:
$\begin{split}\hat{\mathbf{h}}_{\text{stack}}=\begin{bmatrix}\hat{\mathbf{h}}(\mathbf{\bm{\zeta}}_{1})\\ \vdots\\ \hat{\mathbf{h}}(\mathbf{\bm{\zeta}}_{M})\end{bmatrix}=\begin{bmatrix}\mathbf{A}(\mathbf{\bm{\zeta}}_{1})\\ \vdots\\ \mathbf{A}(\mathbf{\bm{\zeta}}_{M})\end{bmatrix}\mathbf{x}+\mathbf{n}=\mathbf{\Phi}\mathbf{x}+\mathbf{n}.\end{split}$ (26)
This is a standard compressed sensing problem, which can be solved using various sparse recovery algorithms. Orthogonal matching pursuit (OMP) is an efficient greedy algorithm for this sparse channel estimation task [38], which iteratively identifies the support set of the sparse vector $\mathbf{x}$, i.e., the indices of the non-zero entries. In Algorithm 1, the OMP-based sparse channel reconstruction framework is presented for FIM systems. Let $\hat{\mathbf{x}}$ be the recovered sparse parameter vector, the channel for any continuous shape $\mathbf{\bm{\zeta}}_{\text{target}}$ can be synthesized as $\hat{\mathbf{h}}_{\text{OMP}}(\mathbf{\bm{\zeta}}_{\text{target}})=\mathbf{A}(\mathbf{\bm{\zeta}}_{\text{target}})\hat{\mathbf{x}}$. This approach elegantly solves both the low-overhead and continuous-space estimation challenges by moving from direct channel interpolation to physical model reconstruction.
Algorithm 1 OMP-Based Parametric Channel Estimation
0: Observed channel vector $\hat{\mathbf{h}}_{\text{stack}}$, Dictionary matrix $\mathbf{\Phi}$,   Sparsity level $S=L\times G$.
[…]
1: Initialize residual: $\mathbf{r}_{0}\leftarrow\hat{\mathbf{h}}_{\text{stack}}$
2: Initialize support set: $\mathcal{S}_{0}\leftarrow\emptyset$
3:for $s=1$ to $S$ do
4:  Compute correlation: $c_{j}\leftarrow|\mathbf{\Phi}_{j}^{H}\mathbf{r}_{s-1}|,\forall j$
5:  Select index: $j_{s}\leftarrow\arg\max_{j}c_{j}$
6:  Update support: $\mathcal{S}_{s}\leftarrow\mathcal{S}_{s-1}\cup\{j_{s}\}$
7:  Solve least squares: $\mathbf{x}_{\mathcal{S}_{s}}\leftarrow(\mathbf{\Phi}_{\mathcal{S}_{s}})^{-1}\hat{\mathbf{h}}_{\text{stack}}$
8:  Update residual: $\mathbf{r}_{s}\leftarrow\hat{\mathbf{h}}_{\text{stack}}-\mathbf{\Phi}_{\mathcal{S}_{s}}\mathbf{x}_{\mathcal{S}_{s}}$
9:end for
10: Construct final solution: $\hat{\mathbf{x}}\leftarrow\begin{cases}\mathbf{x}_{\mathcal{S}_{S}}[i],&\text{if }j\in\mathcal{S}_{S},\\ 0,&\text{otherwise}.\end{cases}$
11:return $\hat{\mathbf{x}}$
[…]
@@
[…]
@@
$\quad\text{s.t.}\ -\zeta\leq[\bm{\zeta}_{m}]_{n}\leq\zeta,\ \forall m,n,$
where $\bar{\zeta}$ represents the maximum possible displacement that the hardware actuators can provide without entering a non-linear and unpredictable response regime.
To find a high-quality solution to the coherence minimization problem without exhaustive search, a constructive greedy algorithm can be designed that builds the set of optimal pilot shapes sequentially. At each step, it selects a new shape from a large pool of candidates that, when added to the already selected set, results in the minimal possible mutual coherence for the newly formed dictionary. This approach ensures that each added pilot shape is maximally complementary to the existing ones in terms of distinguishing the dictionary atoms. The detailed procedure is presented in Algorithm 2.
Algorithm 2 Coherence-Minimal Pilot Selection Optimization
@@
[…]
3:for m=2m=2ătoăMM do
[…]
6:end for
[…]
Remark 1: While an iterative greedy search can find a pilot set optimized for minimal coherence, it is often computationally prohibitive. Therefore, we develop a computationally efficient yet highly effective heuristic using the 2D Fourier basis from Eq. (15) to generate the pilot shapes. The effectiveness of this approach is explained by the inner product that defines mutual coherence $\bm{\phi}_{i}^{H}\bm{\phi}_{j}=\sum_{m=1}^{M}\mathbf{a}(\bm{\zeta}_{m},\theta_{i},\phi_{i})^{H}\mathbf{a}(\bm{\zeta}_{m},\theta_{j},\phi_{j})$. This summation serves as a discrete approximation of an integral over the deformation space. When the pilot shapes $\{\bm{\zeta}_{m}\}$ are chosen from an orthogonal basis like the Fourier basis, they act as orthogonal weighting functions in this sum. According to Fourier analysis [33], this process is analogous to projecting one function onto another using an orthogonal basis, which drives the resulting inner product towards zero for non-identical functions. This systematically decorrelates the dictionary atoms and directly minimizes the mutual coherence.
While the OMP framework provides a computationally efficient approach to sparse recovery, its performance is sensitive to noise and critically depends on the true channel paths aligning with the predefined angular grid, a problem known as basis mismatch. To establish a more robust sparse baseline, we frame the channel estimation problem within a sparse Bayesian learning (SBL) framework. SBL overcomes the limitation of OMP by treating the sparse channel vector $\mathbf{x}$ as a random variable and learning its statistical properties directly from the measurement data [39]. The SBL framework recalls the same linear model as in Eq. (28), i.e., $\hat{\mathbf{h}}_{\text{stack}}=\mathbf{\Phi}\mathbf{x}+\mathbf{n}$, we define the likelihood function as
$p(\hat{\mathbf{h}}_{\text{stack}}\mid\mathbf{x};\sigma^{2})=(\pi\sigma^{2})^{-MN}e^{\left(-\frac{1}{\sigma^{2}}\|\hat{\mathbf{h}}_{\text{stack}}-\mathbf{\Phi}\mathbf{x}\|_{2}^{2}\right)}.$ (29)
To promote sparsity, SBL places a hierarchical prior on $\mathbf{x}$. Specifically, each element $x_{d}$ of $\mathbf{x}$ is assumed to be drawn from a zero-mean Gaussian distribution, with its own distinct variance $\alpha_{d}$, which can be expressed as
$\begin{split}p(\mathbf{x}\mid\bm{\alpha})&=\prod_{d=1}^{D}\mathcal{CN}(x_{d}\mid 0,\alpha_{d})\\ &=(\pi)^{-D}\left(\prod_{d=1}^{D}\alpha_{d}\right)^{-1}e^{\left(-\mathbf{x}^{H}\mathcal{A}^{-1}\mathbf{x}\right)},\end{split}$ (30)
where $\bm{\alpha}=[\alpha_{1},\ldots,\alpha_{D}]^{T}$ is a vector of hyperparameters, and $\mathcal{A}=\text{diag}(\bm{\alpha})$. The sparsity is achieved because if a hyperparameter $\alpha_{d}$ is driven to zero during the learning process, the posterior probability of the corresponding coefficient $x_{d}$ will be sharply peaked at zero.
The SBL algorithm iteratively estimates the posterior distribution of $\mathbf{x}$ while simultaneously learning the hyperparameters $\bm{\alpha}$ and the noise variance $\sigma^{2}$ by maximizing the marginal likelihood $p(\hat{\mathbf{h}}_{\text{stack}}\mid\bm{\alpha},\sigma^{2})$. A common approach to solve this is using an expectation-maximization algorithm. The posterior distribution of $\mathbf{x}$ given the hyperparameters is Gaussian, and the corresponding mean $\bm{\mu}_{x}$ and covariance $\bm{\Sigma}_{x}$ of $\mathbf{x}$ can be expressed as
$\bm{\mu}_{x}=\frac{1}{\sigma^{2}}\bm{\Sigma}_{x}\mathbf{\Phi}^{H}\hat{\mathbf{h}}_{\text{stack}}.$ (31)
$\bm{\Sigma}_{x}=\left(\frac{1}{\sigma^{2}}\mathbf{\Phi}^{H}\mathbf{\Phi}+\mathcal{A}^{-1}\right)^{-1}.$ (32)
The posterior mean $\bm{\mu}_{x}$ serves as the estimate for the sparse vector $\mathbf{x}$. In each iteration, after computing the posterior statistics, the hyperparameters are updated using the following rules [39]:
$\alpha_{d}^{\text{new}}=\left|(\bm{\mu}_{x})_{d}\right|^{2}+(\bm{\Sigma}_{x})_{d,d}\quad\text{for }d=1,\ldots,D,$ (33)
$(\sigma^{2})^{\text{new}}=\frac{\|\hat{\mathbf{h}}_{\text{stack}}-\mathbf{\Phi}\bm{\mu}_{x}\|_{2}^{2}+\text{Re}\left\{\text{Tr}\left(\mathbf{\Phi}\bm{\Sigma}_{x}\mathbf{\Phi}^{H}\right)\right\}}{MN}.$ (34)
Upon convergence, many of the values in $\bm{\alpha}$ will approach zero. We take the final posterior mean $\hat{\mathbf{x}}=\bm{\mu}_{x}$ as the recovered sparse vector. The channel for any continuous target shape $\zeta_{\text{target}}$ can then be synthesized in the same way as OMP, i.e., $\hat{\mathbf{h}}_{\text{SBL}}(\mathbf{\bm{\zeta}}_{\text{target}})=\mathbf{A}(\mathbf{\bm{\zeta}}_{\text{target}})\hat{\mathbf{x}}$. By learning the priors from the data, SBL provides a more principled and robust solution to the sparse recovery problem, making it a strong baseline for comparison.
@@
The above conventional model-based channel estimation frameworks, while foundational, exhibit certain inherent limitations when confronted with the high complexity of the FIM channel estimation problem. The interpolation-based approach, provides the stable linear estimate, yet the fundamental relationship between the FIM deformation vector $\bm{\zeta}$ and the resulting channel vector $\mathbf{h}(\bm{\zeta})$ is highly non-linear. Consequently, the linear estimator or kernel-based models are inherently suboptimal and cannot fully capture the intricate channel dynamics. On the other hand, sparsity-based parametric methods based on sparse recovery are constrained by rigid assumptions that may not hold in practice. Their efficacy critically hinges on two conditions: 1) the channel must be sufficiently sparse, which is often not the case in rich scattering environments; and 2) the true path angles must align with a predefined discrete grid $\mathbf{A}(\bm{\zeta})$ in parametric methods, making the methods highly susceptible to basis mismatch errors. In particular, the true number of paths $S=L\times G$ is unknown to the estimators in practical communication systems, while the performance of parametric methods is critically dependent on this value. These limitations motivate the development of a more flexible and data-driven approach.
@@
IV Learning-Based Channel Estimation with Fourier Neural Operator for FIM Systems
@@
To overcome the aforementioned limitations of model-based channel estimation approaches, in this section, we develop a learning-based framework that moves beyond explicit model assumptions of linearity or sparsity. Specifically, we first provide a theoretical justification for learning-based channel estimation from a Bayesian inference perspective. Then, we reframe the channel estimation problem as one of learning a continuous and non-linear operator that maps any given FIM deformation shape to its corresponding channel response. For this purpose, we propose an H-FNO architecture uniquely suited for this task due to its inherent mesh-independence and its design for learning solutions to physical systems governed by partial differential equations (PDEs). In fact, the propagation of electromagnetic waves is governed by Maxwell’s equations, a system of PDEs. The resulting wireless channel can thus be viewed as the solution to the underlying wave equation, where the geometric shape of the FIM, parameterized by the deformation $\mathbf{\bm{\zeta}}$, acts as a critical boundary condition. Consequently, the task of estimating the channel map ${\mathbf{h}}(\mathbf{\bm{\zeta}})$ is equivalent to learning the solution operator of this governing PDE. The FNO architecture is specifically designed to approximate such operators that map between infinite-dimensional function spaces, making it a natural and powerful framework for this physics-based problem.
@@ IV-A FIM Channel Estimation from a Bayesian Perspective
In essence, channel estimation can be framed as a Bayesian problem of inferring the most plausible channel function $\bar{\mathbf{h}}$, given a set of noisy measurements, $\{\hat{\mathbf{h}}(\bm{\zeta}_{m})\}_{m=1}^{M}$. According to Bayes’ theorem, the posterior probability distribution of the channel function is given by [40]
$p\left(\bar{\mathbf{h}}\mid\{\hat{\mathbf{h}}(\bm{\zeta}_{m})\}\right)\propto p\left(\{\hat{\mathbf{h}}(\bm{\zeta}_{m})\}\mid\bar{\mathbf{h}}\right)p(\bar{\mathbf{h}}),$ (35)
where the likelihood term $p(\{\hat{\mathbf{h}}(\bm{\zeta}_{m})\}\mid\bar{\mathbf{h}})$ is determined by the noise model.
@@
The goal of channel estimation is to find $\bar{\mathbf{h}}$, which corresponds to the Maximum A Posteriori (MAP) estimate [40]. The MAP estimator maximizes the posterior probability, which is equivalent to minimizing the negative log-posterior:
$\bar{\mathbf{h}}_{\text{MAP}}=\arg\min_{\bar{\mathbf{h}}}\left(-\log p\left(\{\hat{\mathbf{h}}(\bm{\zeta}_{m})\}\mid\bar{\mathbf{h}}\right)-\log p(\bar{\mathbf{h}})\right).$ (36)
Given the AWGN model in Eq. (12), the likelihood function is given by
$p(\{\hat{\mathbf{h}}(\bm{\zeta}_{m})\}\mid\bar{\mathbf{h}})=\prod_{m=1}^{M}\frac{1}{(\pi\sigma_{w}^{2})^{N}}e^{-\frac{\|\hat{\mathbf{h}}(\bm{\zeta}_{m})-\bar{\mathbf{h}}(\bm{\zeta}_{m})\|_{2}^{2}}{\sigma_{w}^{2}}}.$ (37)
Taking the negative logarithm of the likelihood gives
$\log p(\{\hat{\mathbf{h}}(\bm{\zeta}_{m})\}\mid\bar{\mathbf{h}})=\sum_{m=1}^{M}\frac{\|\hat{\mathbf{h}}(\bm{\zeta}_{m})-\bar{\mathbf{h}}(\bm{\zeta}_{m})\|_{2}^{2}}{\sigma_{w}^{2}}+C,$ (38)
where $C=MN\log(\pi\sigma_{w}^{2})$ is a constant independent of $\bar{\mathbf{h}}$. Substituting this back into the MAP minimization problem, we get
$\bar{\mathbf{h}}_{\text{MAP}}=\arg\min_{\bar{\mathbf{h}}}\left(\sum_{m=1}^{M}\frac{\|\hat{\mathbf{h}}(\bm{\zeta}_{m})-\bar{\mathbf{h}}(\bm{\zeta}_{m})\|_{2}^{2}}{\sigma_{w}^{2}}-\log p(\bar{\mathbf{h}})\right).$ (39)
Multiplying the objective function by the positive constant $\frac{\sigma_{w}^{2}}{2}$ does not change the minimizer $\bar{\mathbf{h}}$. This transforms the MAP estimation problem into the canonical form of a regularized optimization problem:
@@
$\begin{split}\bar{\mathbf{h}}_{\text{MAP}}=\arg\min_{\bar{\mathbf{h}}}\left(\sum_{m=1}^{M}\frac{1}{2}\|\hat{\mathbf{h}}(\bm{\zeta}_{m})-\bar{\mathbf{h}}(\bm{\zeta}_{m})\|_{2}^{2}-\frac{\sigma_{w}^{2}}{2}\log p(\bar{\mathbf{h}})\right),\end{split}$ (40)
where the squared $\ell_{2}$-norm term acts as a data fidelity term.
@@
This final expression provides that any channel estimator can be interpreted as a solver for this problem. The first term enforces fidelity to the measurements, while the second term $2\sigma_{w}^{2}\log p(\bar{\mathbf{h}})$, acts as a regularizer that is fundamentally defined by the choice of the prior $p(\bar{\mathbf{h}})$. Consequently, the choice of the estimator is fundamentally governed by the choice of the prior $p(\bar{\mathbf{h}})$, which represents our knowledge or assumptions about the channel structure. The model-based methods in Section III make explicit, but often restrictive, prior assumptions, e.g., a strict sparsity prior for OMP estimator. To motivate a more general and physically-grounded prior that can serve as the foundation for a learning-based approach, we now analyze the inherent spectral properties of the FIM channel model itself. As defined in Eq. (4), the FIM channel is a linear superposition of $S=L\times G$ steering vectors, each corresponding to a distinct path angle $(\theta_{s},\phi_{s})$. The response of the $n$-th antenna to the $s$-th path can be expressed as a complex exponential of its spatial coordinates $(x_{n},z_{n})$:
$\mathbf{a}(\bm{\zeta},\theta_{s},\phi_{s})]_{n}=e^{j\left(k_{x,s}x_{n}+k_{z,s}z_{n}+\frac{2\pi}{\lambda}\zeta_{n}\sin\theta_{s}\sin\phi_{s}\right)},$ (41)
@@
where $\alpha_{s}^{\prime}(\bm{\zeta})$ is the effective complex gain of the $s$-th path, which incorporates the original path gain $\alpha_{s}$ and the spatially-varying phase shift induced by the FIM deformation.
Since the Fourier transform of a complex sinusoid is a Dirac delta function, the spatial spectrum $\tilde{\mathbf{h}}(\bm{\zeta};\omega_{x},\omega_{z})$ of the channel is non-zero only at a discrete set of $S$ spatial frequencies. This derivation proves a fundamental property: for any given deformation $\bm{\zeta}$, the FIM channel is inherently sparse in the spatial frequency domain, which provides a rigorous physical justification for selecting an superior prior $p(\bar{\mathbf{h}})$. A function that is sparse in the frequency domain is a specific instance of a broader class of functions whose Fourier coefficients decay rapidly. Consequently, a powerful and general prior for this problem should favor functions with this spectral decay characteristic. By operating directly in the Fourier domain and inherently focusing on a limited number of modes [41], the FNO architecture is naturally aligned with the physical properties of the FIM channel.